Weighted Average Grade.


Suppose that your grade for the class is dependent upon four marking


periods (MP1, MP2, MP3, MP4), each worth 20%, a midterm (MT) worth 10%, and a final exam (FE)


worth 10%. Below is a table of grades for the entire year. Calculate the weighted average grade

to obtain your final average.


Grade/Term


MP1


MP2


MT


MP3


MP4


FE


Grade


84


86


75


94


95


80




1
Expert's answer
2020-10-25T18:46:51-0400

in the following table given grade and percentage written in decimal .(divide percetnage by 100 to convert it to decimal):



formula: Weighted average= "\frac{\sum(grades*frequency)}{\sum (frequency)}"

="\frac{84.7}{1.0}"

=84.7